# Optimizing Layout of Recursive Datatypes with Marmoset

### v2

@@ 5.3 Binary Tree benchmarks
We evaluate Marmoset on a few binary tree benchmarks. These include adding one to all values stored in the internal nodes of the tree, exponentiation on integers stored in internal nodes, copying a tree and getting the right-most leaf value in the tree. For the first three benchmarks, the tree representation we use is:
@@
5.3.1 Add one tree
This benchmark takes a full binary tree and increments the values stored in the internal nodes of the tree. We show the performance of an aligned preorder, inorder and postorder traversal in addition to a misaligned preorder and postorder traversal of the tree. Aligned traversals are ones where the data representation exactly matches the traversal order, for instance, a preorder traversal on a preorder representation of the tree. A misaligned traversal order is where the access patterns of the traversal don’t match the data layout of the tree. For instance, a postorder traversal on a tree serialized in preorder. Table 2 shows the performance numbers. Msolver picks the aligned postorder traversal order which is best performing. It makes the recursive calls to the left and right children of the tree first and increments the values stored in the internal nodes once the recursive calls return. The tree representation is also changed to a postorder representation with the Int placed after the left and right children of the tree. This is in part due to the structure of arrays effect, as the Int are placed closer to each other. Mgreedy on the other hand picks the aligned preorder traversal because of its greedy strategy which prioritizes placing the Int before the left and right subtree. The tree depth was set to 27. At this input size, the Misalignedpost traversal failed due to memory errors. We see significant slowdowns with Misalignedpre, $\sim$35$\times$ slower, because of the skewed access patterns of the traversal.
Table 2: Binary Tree benchmarks runtime statistics in seconds (s). Misalignedpre– A post order traversal on a pre order layout of the tree. Misalignedpost– A pre order traversal on a post order layout of the tree. Alignedpre– A pre order traversal on a pre order layout of the tree. Alignedin– An in order traversal on an in order layout of the tree. Alignedpost– A post order traversal on a post order layout of the tree.
Gibbon Marmoset
Misalignedpre Misalignedpost Alignedpre Alignedin Alignedpost Mgreedy Msolver
@@ 5.3.1 Add one tree
Mean 45.506 memory error 1.318 870 740 740 740 71.31887074074074071.318\,870\,740\,740\,740\,71.318 870 740 740 740 7 1.327 467 398 148 148 41.32746739814814841.327\,467\,398\,148\,148\,41.327 467 398 148 148 4 1.290 646 222 222 222 41.29064622222222241.290\,646\,222\,222\,222\,41.290 646 222 222 222 4 1.318 915 222 222 222 21.31891522222222221.318\,915\,222\,222\,222\,21.318 915 222 222 222 2 1.278 920 282 828 282 41.27892028282828241.278\,920\,282\,828\,282\,41.278 920 282 828 282 4
Median 45.507 memory error 1.291 547 51.29154751.291\,547\,51.291 547 5 1.301 8031.3018031.301\,8031.301 803 1.288 3071.2883071.288\,3071.288 307 1.291 284 51.29128451.291\,284\,51.291 284 5 1.275 4091.2754091.275\,4091.275 409
ub 45.661 memory error 1.334 508 271 008 3761.3345082710083761.334\,508\,271\,008\,3761.334 508 271 008 376 1.342 853 017 459 5031.3428530174595031.342\,853\,017\,459\,5031.342 853 017 459 503 1.296 121 611 091 341 61.29612161109134161.296\,121\,611\,091\,341\,61.296 121 611 091 341 6 1.334 694 924 488 772 71.33469492448877271.334\,694\,924\,488\,772\,71.334 694 924 488 772 7 1.284 846 871 407 015 61.28484687140701561.284\,846\,871\,407\,015\,61.284 846 871 407 015 6
lb 45.351 memory error 1.303 233 210 473 105 51.30323321047310551.303\,233\,210\,473\,105\,51.303 233 210 473 105 5 1.312 081 778 836 793 81.31208177883679381.312\,081\,778\,836\,793\,81.312 081 778 836 793 8 1.285 170 833 353 103 11.28517083335310311.285\,170\,833\,353\,103\,11.285 170 833 353 103 1 1.303 135 519 955 671 81.30313551995567181.303\,135\,519\,955\,671\,81.303 135 519 955 671 8 1.272 993 694 249 549 21.27299369424954921.272\,993\,694\,249\,549\,21.272 993 694 249 549 2
Exponentiation Tree
Mean 45.520 memory error 1.338 833 731 481 481 81.33883373148148181.338\,833\,731\,481\,481\,81.338 833 731 481 481 8 1.340 895 601 851 8521.3408956018518521.340\,895\,601\,851\,8521.340 895 601 851 852 1.296 339 515 151 515 11.29633951515151511.296\,339\,515\,151\,515\,11.296 339 515 151 515 1 1.336 237 537 037 037 21.33623753703703721.336\,237\,537\,037\,037\,21.336 237 537 037 037 2 1.294 804 525 252 5251.2948045252525251.294\,804\,525\,252\,5251.294 804 525 252 525
Median 45.567 memory error 1.312 967 51.31296751.312\,967\,51.312 967 5 1.314 778 51.31477851.314\,778\,51.314 778 5 1.293 3761.2933761.293\,3761.293 376 1.310 2921.3102921.310\,2921.310 292 1.291 8291.2918291.291\,8291.291 829
ub 45.702 memory error 1.354 349 493 420 148 81.35434949342014881.354\,349\,493\,420\,148\,81.354 349 493 420 148 8 1.356 565 502 024 6421.3565655020246421.356\,565\,502\,024\,6421.356 565 502 024 642 1.301 961 583 230 177 41.30196158323017741.301\,961\,583\,230\,177\,41.301 961 583 230 177 4 1.352 130 068 835 353 51.35213006883535351.352\,130\,068\,835\,353\,51.352 130 068 835 353 5 1.300 631 257 753 884 31.30063125775388431.300\,631\,257\,753\,884\,31.300 631 257 753 884 3
lb 45.338 memory error 1.323 317 969 542 814 81.32331796954281481.323\,317\,969\,542\,814\,81.323 317 969 542 814 8 1.325 225 701 679 0621.3252257016790621.325\,225\,701\,679\,0621.325 225 701 679 062 1.290 717 447 072 852 91.29071744707285291.290\,717\,447\,072\,852\,91.290 717 447 072 852 9 1.320 345 005 238 7211.3203450052387211.320\,345\,005\,238\,7211.320 345 005 238 721 1.288 977 792 751 165 91.28897779275116591.288\,977\,792\,751\,165\,91.288 977 792 751 165 9
@@
Mean 45.524 memory error 1.312 177 046 296 2961.3121770462962961.312\,177\,046\,296\,2961.312 177 046 296 296 1.323 082 518 518 518 71.32308251851851871.323\,082\,518\,518\,518\,71.323 082 518 518 518 7 1.281 644 191 919 191 91.28164419191919191.281\,644\,191\,919\,191\,91.281 644 191 919 191 9 1.311 649 712 962 962 71.31164971296296271.311\,649\,712\,962\,962\,71.311 649 712 962 962 7 1.278 667 242 424 242 51.27866724242424251.278\,667\,242\,424\,242\,51.278 667 242 424 242 5
Median 45.529 memory error 1.286 019 51.28601951.286\,019\,51.286 019 5 1.296 986 51.29698651.296\,986\,51.296 986 5 1.278 1531.2781531.278\,1531.278 153 1.284 7831.2847831.284\,7831.284 783 1.276 031.276031.276\,031.276 03
ub 45.673 memory error 1.327 728 808 958 8631.3277288089588631.327\,728\,808\,958\,8631.327 728 808 958 863 1.338 750 685 128 811.338750685128811.338\,750\,685\,128\,811.338 750 685 128 81 1.287 317 967 848 180 51.28731796784818051.287\,317\,967\,848\,180\,51.287 317 967 848 180 5 1.327 481 317 169 239 71.32748131716923971.327\,481\,317\,169\,239\,71.327 481 317 169 239 7 1.284 358 794 308 634 31.28435879430863431.284\,358\,794\,308\,634\,31.284 358 794 308 634 3
lb 45.374 memory error 1.296 625 283 633 7291.2966252836337291.296\,625\,283\,633\,7291.296 625 283 633 729 1.307 414 351 908 227 31.30741435190822731.307\,414\,351\,908\,227\,31.307 414 351 908 227 3 1.275 970 415 990 203 21.27597041599020321.275\,970\,415\,990\,203\,21.275 970 415 990 203 2 1.295 818 108 756 685 81.29581810875668581.295\,818\,108\,756\,685\,81.295 818 108 756 685 8 1.272 975 690 539 850 61.27297569053985061.272\,975\,690\,539\,850\,61.272 975 690 539 850 6
5.3.2 Exponentiation Tree
This traversal does exponentiation on the values stored in the internal nodes of the tree. It is more computationally intensive than incrementing the value. We raise the Int to a power of 10 on a tree of depth 27. Table 2 shows the performance of the different layout and traversal orders. Msolver picks the Alignedpost representation which is the best performing, whereas Mgreedy picks the Alignedpre representation.
5.3.3 Copy Tree
Copy-tree takes a full binary tree and makes a fresh copy of the tree in a new memory location. We use a tree of depth 27 in our evaluation. Table 2 shows the performance of different layout and traversal orders. We see that Alignedpost traversal performs the best. Indeed, Msolver picks the Alignedpost representation, whereas, Mgreedy chooses the Alignedpre representation.
5.3.4 Rightmost
This traversal does recursion on the right child of the tree and returns the Int value stored in the right-most leaf of the tree. Figure 8 shows an example of a tree with two different serializations of the tree: left-to-right (top) and right-to-left (bottom). The right-to-left serialization is more efficient because the constant-step movements (blue arrows) are usually more favorable than variable-step ones (red arrows) on modern hardware. Both Msolver and Mgreedy pick the right-to-left serialization, and Table 1(c) shows that this choice performs better in the benchmark.
@@ 5.4 Blog Software Case Study
The Blog software case study serves as an example of a realistic benchmark, representing a sample of components from a blog management web service. The main data structure is a linked list of blogs where each blog contains fields such as header, ID, author information, content, hashtags and date. The fields are a mix of recursive and non-recursive datatypes. For instance, Content is a recursive type (the Pandoc Block type), but Author is a single string wrapped in a data constructor. One possible permutation of fields in the blog is:
@@
We evaluate Marmoset’s performance using three different traversals over a list of blogs. Overall, these traversals only use three fields: Content and Tags, and the tail of the linked-list, Blogs. The fields used by each individual traversal are referred as active fields and the rest are referred as passive fields, and we specify these per-traversal in the following. In Table 3 we report the performance of all six layouts obtained by permuting the order of these three fields, and two additional layouts (Columns 1 and 2). The column names are indicative of the order of fields used. For example, the column hiadctb reports numbers for a layout where the fields are ordered as: Header, Id, Author, Date, Content, HashTags, and Blogs (the tail of the list). We use Gibbon to gather these run times. The second to last and the last two columns show the run times of these traversals using Marmoset’s greedy and solver based optimization. The layout it picks is tailored to each individual traversal, so we specify it per-traversal next.
Table 3: Runtime statistics in seconds (s) for different blog traversals for a subset permutation of different layouts when compiled with Gibbon. Legend: h - Header, t - HashTags, b - Blogs, i - TagID, c - Content, a - Author, d - Date.
Gibbon Marmoset
@@
Mean 2.032 400 527 777 7782.0324005277777782.032\,400\,527\,777\,7782.032 400 527 777 778 2.017 655 231 481 4812.0176552314814812.017\,655\,231\,481\,4812.017 655 231 481 481 2.768 484 333 333 333 32.76848433333333332.768\,484\,333\,333\,333\,32.768 484 333 333 333 3 1.716 654 833 333 333 41.71665483333333341.716\,654\,833\,333\,333\,41.716 654 833 333 333 4 2.802 108 222 222 $2$.8021082222222222.802\,108\,222\,222\,$2$.802 108 222 222 222 2.809 880 222 222 $2$.8098802222222222.809\,880\,222\,222\,$2$.809 880 222 222 222 1.857 133 740 740 740 71.85713374074074071.857\,133\,740\,740\,740\,71.857 133 740 740 740 7 1.801 490 703 703 703 31.80149070370370331.801\,490\,703\,703\,703\,31.801 490 703 703 703 3 1.733
Median 1.993 221.993221.993\,221.993 22 1.976 430 51.97643051.976\,430\,51.976 430 5 2.766 2692.7662692.766\,2692.766 269 1.661 7591.6617591.661\,7591.661 759 2.801 9012.8019012.801\,9012.801 901 2.809 4932.8094932.809\,4932.809 493 1.804 6841.8046841.804\,6841.804 684 1.746 745 499 999 999 81.74674549999999981.746\,745\,499\,999\,999\,81.746 745 499 999 999 8 1.724
ub 2.056 570 170 215 6742.0565701702156742.056\,570\,170\,215\,6742.056 570 170 215 674 2.042 352 241 875 830 62.04235224187583062.042\,352\,241\,875\,830\,62.042 352 241 875 830 6 2.773 769 727 132 014 82.77376972713201482.773\,769\,727\,132\,014\,82.773 769 727 132 014 8 1.749 551 172 566 805 61.74955117256680561.749\,551\,172\,566\,805\,61.749 551 172 566 805 6 2.805 736 222 347 7742.8057362223477742.805\,736\,222\,347\,7742.805 736 222 347 774 2.811 850 435 449 460 72.81185043544946072.811\,850\,435\,449\,460\,72.811 850 435 449 460 7 1.888 528 756 113 495 41.88852875611349541.888\,528\,756\,113\,495\,41.888 528 756 113 495 4 1.834 483 954 542 924 61.83448395454292461.834\,483\,954\,542\,924\,61.834 483 954 542 924 6 1.746 409 750 718 698 21.74640975071869821.746\,409\,750\,718\,698\,21.746 409 750 718 698 2
lb 2.008 230 885 339 8822.0082308853398822.008\,230\,885\,339\,8822.008 230 885 339 882 1.992 958 221 087 131 51.99295822108713151.992\,958\,221\,087\,131\,51.992 958 221 087 131 5 2.763 198 939 534 6522.7631989395346522.763\,198\,939\,534\,6522.763 198 939 534 652 1.683 758 494 099 861 11.68375849409986111.683\,758\,494\,099\,861\,11.683 758 494 099 861 1 2.798 480 222 096 672.798480222096672.798\,480\,222\,096\,672.798 480 222 096 67 2.807 910 008 994 983 62.80791000899498362.807\,910\,008\,994\,983\,62.807 910 008 994 983 6 1.768 497 452 864 482 11.76849745286448211.768\,497\,452\,864\,482\,11.768 497 452 864 482 1 1.719 111 522 008 5741.7191115220085741.719\,111\,522\,008\,5741.719 111 522 008 574 1.633 349 537 273 928 61.63334953727392861.633\,349\,537\,273\,928\,61.633 349 537 273 928 6
5.4.1 FilterBlogs
This traversal filters the list of blogs and only retains those which contain the given hash-tag in their HashTags field. Since Blogs, Content, and HashTags are recursive fields, changing their layout should represent greater differences in performance. The active fields for this traveral are HashTags and Blogs. Theoretically, the performance of this traversal is optimized when the HashTags field is serialized before Blogs on account of the first access to HashTags in the traversal. This is confirmed in practice with the layout tbchiad being the fastest. The access to HashTags comes first since its serialized first in memory and once HashTags is fully traversed, the traversal can access blogs in an optimal manner without unnecessary movement in the packed buffer.
@@ 5.4.1 FilterBlogs
Indeed, the layout chosen by Marmoset puts HashTags first followed by Blogs. The order of the rest of the fields remains unchanged compared to the source program, but this has no effect on performance since they are passive fields. As shown in Table 3, the performance of the layout chosen by Marmoset is similar to the layout tbchiad when compiled with Gibbon. Both Msolver and Mgreedy pick the layout tbhiadc when compiled from the initial layout hiadctb.
5.4.2 EmphContent
@@
This traversal takes a keyword and a list of blogs as input. It then searches the content of each blog for the keyword and emphasizes all its occurrences if present. If the keyword isn’t present, the content remains unchanged. The active fields in this traversal are Content and Blogs. Therefore, based on the access pattern (Content accessed before Blogs), the layout with the best performance should place Content first followed by Blogs. Indeed, the layout with the best performance when compiled with Gibbon is cbiadht. The layout that Msolver chooses prioritizes the placement of Blogs before Content. It subsequently changes the traversal to recurse on the blogs first and then emphasize content. The other passive fields are all placed afterwards. The layout chosen by Msolver is bchiadt, whereas the layout chosen by Mgreedy is cbhiadt when compiled from the initial layout hiadctb. Table 3 shows the performance of all the layouts for this traversal. The performance of Mgreedy and Msolver differ because data constructors not shown differ in their layout choices.
5.4.3 TagSearch
@@
This traversal takes a keyword and a list of blogs as input. It looks for the presence of a keyword in the HashTags field. If it is present, it emphasizes this keyword in the Content, and leaves it unchanged otherwise. The layout that gives the best performance when compiled with Gibbon is tcbhiad. This is as expected based on the access patterns of the traversal which access HashTags followed by Content followed by Blogs. Msolver chooses the layout tbchiad which places HashTags, followed by Blogs, followed by Content and changes the traversal to recurse on Blogs first and later emphasize Content in the then branch. On the other hand, the layout chosen by Mgreedy is tcbhiad when compiled from the initial layout hiadctb.

### v3

@@ 5.3 Binary Tree Benchmarks
We evaluate Marmoset on a few binary tree benchmarks: adding one to all values in a tree, exponentiation on integers stored in internal nodes, copying a tree and getting the right-most leaf value in the tree. For the first three benchmarks, the tree representation we use is:
@@
Table 3: Run-time mean and 95% confidence interval (ub, lb) for different layouts and traversal orders in the binary tree benchmarks (seconds). Misalgnpre – post-order traversal on the pre-order layout of the tree. Misalgnpost – pre-order traversal on the post-order layout of the tree. Algnpre – pre-order traversal on the pre-order layout of the tree. Algnin – in-order traversal on an in-order layout of the tree. Algnpost – post-order traversal on the post-order layout of the tree.
Benchmark name Gibbon Marmoset
Misalgnpre Misalgnpost Algnpre Algnin Algnpost Mgreedy Msolver
AddOneTree 45.51 (45.35, 45.66) memory error 1.29 (1.29, 1.29) 1.30 (1.30, 1.30) 1.29 (1.29, 1.29) 1.29 (1.29, 1.29) 1.28 (1.28, 1.28)
ExpTree 45.52 (45.34, 45.70) memory error 1.311 942 478 873 239 71.31194247887323971.311\,942\,478\,873\,239\,71.311 942 478 873 239 7 (1.311 802 107 498 399 91.31180210749839991.311\,802\,107\,498\,399\,91.311 802 107 498 399 9, 1.312 082 850 248 079 51.31208285024807951.312\,082\,850\,248\,079\,51.312 082 850 248 079 5) 1.312 822 464 788 732 51.31282246478873251.312\,822\,464\,788\,732\,51.312 822 464 788 732 5 (1.312 729 018 973 9891.3127290189739891.312\,729\,018\,973\,9891.312 729 018 973 989, 1.312 915 910 603 4761.3129159106034761.312\,915\,910\,603\,4761.312 915 910 603 476) 1.293 828 563 380 281 71.29382856338028171.293\,828\,563\,380\,281\,71.293 828 563 380 281 7 (1.293 738 256 305 614 11.29373825630561411.293\,738\,256\,305\,614\,11.293 738 256 305 614 1, 1.293 918 870 454 949 41.29391887045494941.293\,918\,870\,454\,949\,41.293 918 870 454 949 4) 1.311 794 732 394 366 21.31179473239436621.311\,794\,732\,394\,366\,21.311 794 732 394 366 2 (1.311 661 357 137 6031.3116613571376031.311\,661\,357\,137\,6031.311 661 357 137 603, 1.311 928 107 651 129 31.31192810765112931.311\,928\,107\,651\,129\,31.311 928 107 651 129 3) 1.292 822 605 633 8031.2928226056338031.292\,822\,605\,633\,8031.292 822 605 633 803 (1.291 955 136 039 502 71.29195513603950271.291\,955\,136\,039\,502\,71.291 955 136 039 502 7, 1.293 690 075 228 103 31.29369007522810331.293\,690\,075\,228\,103\,31.293 690 075 228 103 3)
CopyTree 45.52 (45.37, 45.67) memory error 1.287 096 056 338 028 31.28709605633802831.287\,096\,056\,338\,028\,31.287 096 056 338 028 3 (1.286 875 053 013 4911.2868750530134911.286\,875\,053\,013\,4911.286 875 053 013 491, 1.287 317 059 662 565 61.28731705966256561.287\,317\,059\,662\,565\,61.287 317 059 662 565 6) 1.296 957 943 661 971 61.29695794366197161.296\,957\,943\,661\,971\,61.296 957 943 661 971 6 (1.296 793 427 746 940 61.29679342774694061.296\,793\,427\,746\,940\,61.296 793 427 746 940 6, 1.297 122 459 577 002 61.29712245957700261.297\,122\,459\,577\,002\,61.297 122 459 577 002 6) 1.278 674 281 690 140 91.27867428169014091.278\,674\,281\,690\,140\,91.278 674 281 690 140 9 (1.278 593 097 663 421 21.27859309766342121.278\,593\,097\,663\,421\,21.278 593 097 663 421 2, 1.278 755 465 716 860 51.27875546571686051.278\,755\,465\,716\,860\,51.278 755 465 716 860 5) 1.285 843 183 098 591 51.28584318309859151.285\,843\,183\,098\,591\,51.285 843 183 098 591 5 (1.285 758 527 448 210 41.28575852744821041.285\,758\,527\,448\,210\,41.285 758 527 448 210 4, 1.285 927 838 748 972 51.28592783874897251.285\,927\,838\,748\,972\,51.285 927 838 748 972 5) 1.276 624 929 577 464 81.27662492957746481.276\,624\,929\,577\,464\,81.276 624 929 577 464 8 (1.276 501 266 558 847 81.27650126655884781.276\,501\,266\,558\,847\,81.276 501 266 558 847 8, 1.276 748 592 596 081 81.27674859259608181.276\,748\,592\,596\,081\,81.276 748 592 596 081 8)
@@
This benchmark takes a full binary tree and increments the values stored in the internal nodes of the tree. We show the performance of an aligned preorder, inorder and postorder traversal in addition to a misaligned preorder and postorder traversal of the tree. Aligned traversals are ones where the data representation exactly matches the traversal order, for instance, a preorder traversal on a preorder representation of the tree. A misaligned traversal order is where the access patterns of the traversal don’t match the data layout of the tree. For instance, a postorder traversal on a tree serialized in preorder. Table 3 shows the performance numbers. Msolver picks the aligned postorder traversal order which is best performing. It makes the recursive calls to the left and right children of the tree first and increments the values stored in the internal nodes once the recursive calls return. The tree representation is also changed to a postorder representation with the Int placed after the left and right children of the tree. This is in part due to the structure of arrays effect, as the Int are placed closer to each other. Mgreedy on the other hand picks the aligned preorder traversal because of its greedy strategy which prioritizes placing the Int before the left and right subtree. The tree depth is set to 27. At this input size, the Misalgnpost traversal failed due to memory errors, and Misalgnpre runs $\sim$35$\times$ slower than aligned versions because of the skewed access patterns of the traversal.
ExpTree
This traversal does exponentiation on the values stored in the internal nodes of the tree. It is more computationally intensive than incrementing the value. We raise the Int to a power of 10 on a tree of depth 27. Table 3 shows the performance of the different layout and traversal orders. Msolver picks the Algnpost representation which is the best performing, whereas Mgreedy picks the Algnpre representation.
@@
Copy-tree takes a full binary tree and makes a fresh copy of the tree in a new memory location. We use a tree of depth 27 in our evaluation. Table 3 shows the performance of different layout and traversal orders. We see that Algnpost traversal performs the best. Indeed, Msolver picks the Algnpost representation, whereas, Mgreedy chooses the Algnpre representation.
Rightmost
This traversal does recursion on the right child of the tree and returns the Int value stored in the right-most leaf of the tree. Figure 8 shows an example of a tree with two different serializations of the tree: left-to-right (top) and right-to-left (bottom). The right-to-left serialization is more efficient because the constant-step movements (blue arrows) are usually more favorable than variable-step ones (red arrows) on modern hardware. Both Msolver and Mgreedy pick the right-to-left serialization, and Table 2 shows that this choice performs better in the benchmark.
@@ 5.4 Blog Software Case Study
The Blog software case study serves as an example of a realistic benchmark, representing a sample of components from a blog management web service. The main data structure is a linked list of blogs where each blog contains fields such as header, ID, author information, content, hashtags and date. The fields are a mix of recursive and non-recursive datatypes. For instance, Content is a recursive type (the Pandoc Block type), but Author is a single string wrapped in a data constructor. [^9] At times, we have to wrap scalars in data constructors to make them packed fields. Gibbon does not always support mixing scalar and packed fields due to compiler bugs. One possible permutation of fields in the blog is:
@@
We evaluate Marmoset’s performance using three different traversals over a list of blogs. Overall, the traversals accept a keyword and a list of blogs; in the blogs, the traversals inspect either of the three fields: Content, HashTags, and the tail of the linked-list, Blogs. (Since Blogs, Content, and HashTags are recursive fields, changing their layout should represent greater differences in performance.) The fields used by an individual traversal are referred as active fields and the rest are referred as passive fields, and we specify these per-traversal below.
In Table 4, we report the performance of the six possible layouts obtained by permuting the order of the three recursive fields, and two additional layouts (Columns 1 and 2). The column names indicate the order of fields used; for example, the column hiadctb reports numbers for the layout with fields ordered as: Header, Id, Author, Date, Content, HashTags, and Blogs. All run times are gathered with Gibbon, and last two columns show the run times for code compiled using Marmoset’s greedy and solver-based optimization, respectively.
Table 4: Run-time mean and 95% confidence interval (ub, lb) for different layouts in the blog software benchmarks (seconds). Several possible permutations of layout are shown. Layout names abbreviations: h – Header, t – HashTags, b – Blogs, i – TagID, c – Content, a – Author, d – Date.
Bench. name Gibbon Marmoset
@@
1.989 340 112 676 056 11.98934011267605611.989\,340\,112\,676\,056\,11.989 340 112 676 056 1 (1.989 037 693 588 153 71.98903769358815371.989\,037\,693\,588\,153\,71.989 037 693 588 153 7, 1.989 642 531 763 958 61.98964253176395861.989\,642\,531\,763\,958\,61.989 642 531 763 958 6) 1.979 403 225 352 113 21.97940322535211321.979\,403\,225\,352\,113\,21.979 403 225 352 113 2 (1.979 185 046 541 3941.9791850465413941.979\,185\,046\,541\,3941.979 185 046 541 394, 1.979 621 404 162 832 41.97962140416283241.979\,621\,404\,162\,832\,41.979 621 404 162 832 4) 3.294 649 845 070 422 23.29464984507042223.294\,649\,845\,070\,422\,23.294 649 845 070 422 2 (3.292 692 808 020 684 83.29269280802068483.292\,692\,808\,020\,684\,83.292 692 808 020 684 8, 3.296 606 882 120 159 73.29660688212015973.296\,606\,882\,120\,159\,73.296 606 882 120 159 7) 1.680 002 676 056 337 81.68000267605633781.680\,002\,676\,056\,337\,81.680 002 676 056 337 8 (1.679 670 088 341 0371.6796700883410371.679\,670\,088\,341\,0371.679 670 088 341 037, 1.680 335 263 771 638 71.68033526377163871.680\,335\,263\,771\,638\,71.680 335 263 771 638 7) 3.309 097 267 605 633 63.30909726760563363.309\,097\,267\,605\,633\,63.309 097 267 605 633 6 (3.308 674 578 901 5363.3086745789015363.308\,674\,578\,901\,5363.308 674 578 901 536, 3.309 519 956 309 731 33.30951995630973133.309\,519\,956\,309\,731\,33.309 519 956 309 731 3) 3.302 885 873 239 436 63.30288587323943663.302\,885\,873\,239\,436\,63.302 885 873 239 436 6 (3.302 365 980 422 7843.3023659804227843.302\,365\,980\,422\,7843.302 365 980 422 784, 3.303 405 766 056 089 43.30340576605608943.303\,405\,766\,056\,089\,43.303 405 766 056 089 4) 1.817 019 549 295 774 61.81701954929577461.817\,019\,549\,295\,774\,61.817 019 549 295 774 6 (1.816 491 676 802 975 31.81649167680297531.816\,491\,676\,802\,975\,31.816 491 676 802 975 3, 1.817 547 421 788 573 91.81754742178857391.817\,547\,421\,788\,573\,91.817 547 421 788 573 9) 1.755 978 478 873 239 31.75597847887323931.755\,978\,478\,873\,239\,31.755 978 478 873 239 3 (1.755 503 921 352 512 51.75550392135251251.755\,503\,921\,352\,512\,51.755 503 921 352 512 5, 1.756 453 036 393 9661.7564530363939661.756\,453\,036\,393\,9661.756 453 036 393 966) 1.736 654 943 661 9721.7366549436619721.736\,654\,943\,661\,9721.736 654 943 661 972 (1.736 380 487 003 4411.7363804870034411.736\,380\,487\,003\,4411.736 380 487 003 441, 1.736 929 400 320 503 11.73692940032050311.736\,929\,400\,320\,503\,11.736 929 400 320 503 1)
FilterBlogs
@@
filters the list of blogs and only retains those which contain the given keyword in the HashTags field. The active fields for this traveral are HashTags and Blogs. Theoretically, the performance of this traversal is optimized when the HashTags field is serialized before Blogs on account of the first access to HashTags in the traversal. This is confirmed in practice with the layout tbchiad being the fastest. Marmoset chooses the layout with HashTags serialized first and followed by Blogs; the order of other fields remains unchanged compared to the source program, but this has no effect on performance since they are passive fields. Table 4 also shows that layout chosen by Marmoset performs similar to the layout tbchiad. Both Msolver and Mgreedy pick the layout tbhiadc when compiled from the initial layout hiadctb.
EmphContent
@@
searches the content of each blog for the keyword and emphasizes all its occurrences there (if any). The active fields in this traversal are Content and Blogs. Based on the access pattern (Content accessed before Blogs), the layout with the best performance should place Content first followed by Blogs. In practice, the layout with the best performance is cbiadht. In contrast, Msolver prioritizes the placement of Blogs before Content, but it also changes the traversal to recurse on the blogs first and then emphasize content. The passive fields are placed afterwards. The layout chosen by Msolver is bchiadt, whereas the layout chosen by Mgreedy is cbhiadt when compiled from the initial layout hiadctb. The performance of Mgreedy and Msolver differ because datatypes other than Blogs differ in their layout choices.
TagSearch
@@
looks for the presence of the keyword in the HashTags field, and if the keyword is present, the traversal emphasizes the keyword in the Content. The layout with the best performance is tcbhiad because of the access pattern, which inspects HashTags followed by Content followed by Blogs. Msolver chooses the layout tbchiad—which places HashTags followed by Blogs followed by Content—and changes the traversal to recurse on Blogs first and later emphasize Content in the then branch. On the other hand, Mgreedy chooses tcbhiad when compiled from the initial layout hiadctb.